OpenStudy (anonymous):

6. Starting from 1.5 miles away, a car drives toward a speed checkpoint and then passes it. The car travels at a constant rate of 53 miles per hour. The distance of the car from the checkpoint is given by d = |1.5 – 53t|. At what times is the car 0.1 miles from the checkpoint? Calculate your answer in seconds. (1 point) 95.1s and 108.7s 10.2s and 101.9s 108.7s and 10.2s 95.1s and 10.2s

OpenStudy (anonymous):

is any damn body going to help mw

OpenStudy (anonymous):

Well, the first thing we want to do is convert miles per hour into miles per second. We do this by the following multiplication:\[\frac{ miles }{ hour } \times \frac{ 1 hour }{ 60 minutes } \times \frac{ 1minute }{ 60 seconds }\]So\[53 \times \frac{ 1 }{ 60 } \times \frac{ 1 }{ 60 } = \frac{ 53 }{ 3600 }\approx 0.014722 \frac{ miles }{ second }\]Now, we know that there must be two solutions to the problem. To simplify the problem we will instead consider the displacement between the checkpoint and the car. (Displacement differs from distance in that displacement gives the distance and direction). This lets us drop the absolute value. So now we have\[D=1.5-0.014722t\]Now, the two values of D that we care about are D=0.1 and D=-0.1 So the two times will just be solutions of the following two equations:\[0.1=1.5-0.014722t\]and\[-0.1=1.5-0.014722t\]

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

no not much that's why i was just looking for the answer itself

OpenStudy (anonymous):

What part is confusing?

OpenStudy (anonymous):

@marymaryquitecontrary I can't just give you the answer, but I can help you get there yourself.